OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if z was isolated in the first equation. 2x – 5y + z = 15 4x – y + 3z = 15 3x – 2y + 2z = 13

OpenStudy (anonymous):

1) z = 15 - 2x + 5y which would result in 2) 4x -y + 3(15 -2x +5y) = 15 4x -y + 45 -6x + 15y = 15 -2x +14y = -30 3) 3x -2y +2(15 - 2x +5y) = 13 3x - 2y + 30 - 4x + 10y = 13 -x +8y = -17

OpenStudy (anonymous):

which is the answer

OpenStudy (anonymous):

these are the answers given to choose from –2x + 14y = –30 –x + 8y = –17 10x – 16y = –30 7x – 12y = –15 –18x – 14z = –60 –5x + 8z = 43 22x + 16z = 90 11x + 8z = 43

OpenStudy (anonymous):

hey can u please help me understand what u just wrote

OpenStudy (anonymous):

there i answered